what is 3(x-4) - 4(x-3) = x+3-(x-2)
Answer by kiru_khandelwal(79)

3(x-4) - 4(x-3) = x+3-(x-2)
=> 3x -12 -4x + 12 = x + 3 -x + 2
=> -x = 5
=> x = -5